OpenStudy (anonymous):

whats one ninth of 45?

OpenStudy (anonymous):

1/9 * 45 = 45/9 = 5?

OpenStudy (anonymous):

yep its 5

OpenStudy (anonymous):

okay thanks! i dont know how to find it

OpenStudy (anonymous):

You divide 45 by 9.

OpenStudy (anonymous):

one ninth = one over nine = 1/9 of stands for multiplication 45 is just a number so 1/9 x 45

OpenStudy (anonymous):

linguistically 'of' means multiply. \[\text{one ninth} = \frac{1}{9}\] So \[\text{one ninth of 45} = \frac{1}{9}\times45\]

OpenStudy (anonymous):

ohh